(2 points) x19 20 21 22 23 P(X = x) 0.1 0.1 0.2 0.1 0.5 Given the discrete probability distribution above, determine the following (a) P(X = 21) = (b) P(X > 20) = (c) P(X> 20)=
Answer #1

Given that, the discrete probability distrinution is as follows,

x 19 20 21 22 23
P(X = x) 0.1 0.1 0.2 0.1 0.5

To determine the following probabilities,

a) P(X = 21) = 0.2

b) P(X > 20) = P(X = 21) + P(X = 22) + P(X = 23)

= 0.2 + 0.1 + 0.5

P(X > 20 ) = 0.8

c) P(X \geq 20) = P(X = 20) + P(X =21)+P(X=22)+P(X=23)

=0.1+ 0.2+0.1+0.5

  P(X \geq 20) = 0.9

OR

P(X \geq 20) = 1 - P(X< 20) = 1 - P(X= 19) = 1-0.1 = 0.9